[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun Jun 11, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Jun 12 03:01:27 UTC 2017


HSU Shift Log for Sun Jun 11 2017 23:01:21 GMT-0400 (EDT)

Shift: R

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: ATCF changed initial status of TD-03 at 18Z and 00Z from "TD"  
to "Low" despite system being initialized as 30-kt as a TD.

N-AWIPS issues: None.

AWIPS-2/NCP issues: None.

Other Techincal issues: None.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170612/b6412b0c/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list